OpenStudy (anonymous):

Lani is seven years older than her brother. Write an expression for the age of Lani eight years ago......Please give me the answer....... Thank You

OpenStudy (anonymous):

let Lani be X and her brother be Y

OpenStudy (anonymous):

so presently.. Y + 7= X

OpenStudy (dan815):

Lanis age = 7+Brothers age

OpenStudy (anonymous):

so eight years ago.. Lani was X- 8 and her brother was Y- 8

OpenStudy (anonymous):

since Lani present age is X and X = Y + 7 Therefore,eight years ago.. Lani will be ( X - 8) = (Y + 7) - 8

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

ya welcome